OpenStudy (anonymous):

Find the equation of an ellipse with center at the origin passing through (-3,2) and its latus rectum is three times the distance from the center .

rebeccaxhawaii (rebeccaxhawaii):

WELCOME TO OPENSTUDY !!!

ILovePuppiesLol (ilovepuppieslol):

0_____________________________________________0 i heard @rebeccaxhawaii is the best at this kinda stuff

rebeccaxhawaii (rebeccaxhawaii):

@Astrophysics

OpenStudy (anonymous):

Thanks. Can you help me solve that problem?

rebeccaxhawaii (rebeccaxhawaii):

@imqwerty

ILovePuppiesLol (ilovepuppieslol):

@Hero

OpenStudy (abdullahsadiqi):

\[\frac{ (x-h)^2 }{ a^2 }+\frac{ (y-k)^2 }{ b^2 }=1 \]